One challenge in using hydrogen as a fuel, rather than fossil fuels for instance, is how to safely and reliably produce and store the highly flammable gas. Chemists have proposed using metal hydrides for this application. Calcium hydride, for instance, has shown some promise. Calcium hydride reacts with water to form calcium hydroxide and hydrogen gas. Determine the mass of calcium hydride required to fill a 10.0 L fuel tank at a pressure of 15.0 atm at 23 oC.
1
Expert's answer
2020-05-01T14:30:56-0400

The reaction

"CaH_2+2H_2O =>Ca(OH)_2 +2H_2\u2191"

v-numbers of moles. M-Molar mass

m-mass m=v*M

law of ideal gas:

vRT=PV

v=PV/RT

T=296K

V=10l

P=15 at

v(H2)=6.18 mole

v(CaH2)=3,09

m=3.09*42=129,78g